# Black Holes and Marchenko-Pastur distribution

### v2

@@ 1 Introduction
with $\lambda_{\pm}=\left(1\pm\sqrt{c}\right)^{2}$. More precisely, $D(\lambda)=\Omega f(\lambda)$ with $c=\Omega\operatorname{e}^{-S}\in(0,\infty)$. The significance of the MP distribution stems from the MP theorem [5], which can be stated as follows. Let $X$ be an $M\times N$ matrix with independent and identically distributed real or complex entries. Then, in the large $N$ limit with $M/N\to c\in(0,\infty)$ fixed, the asymptotic eigenvalue distribution of the $M\times M$ matrix $A=\frac{1}{N}XX^{\dagger}$ (also called the Wishart matrix) is universal and given by the MP distribution. The MP distribution also appears under the name of free Poisson distribution in Free Probability Theory [6]. Since MP’s original proof from 1967, many different proofs have appeared in the mathematical literature (see [7, 8, 9] and references therein). For physicists, a proof using Feynman diagrams [10] seems especially interesting.
@@ 2.3 State complexity
$C(z)=\int\limits_{0}^{\infty}\,\mathrm{d}t\,C(t)\operatorname{e}^{-zt}\qquad(\Re z>0)~{}.$ (2.24)
@@
$C(z)=\sum\limits_{n=1}^{\infty}n\int\,\mathrm{d}\mu(E)\int\,\mathrm{d}\mu(E^{\prime})\frac{P_{n}(E)P_{n}(E^{\prime})}{h_{n}[z+i(E-E^{\prime})]}~{}.$ (2.25)
@@
$C(z)=-i\sum\limits_{n=1}^{\infty}n\int\,\mathrm{d}\mu(E)\frac{P_{n}(E)Q_{n}(-iz+E)}{h_{n}}~{}.$ (2.26)
@@
$zC(z)=\sum\limits_{n=0}^{\infty}\int\,\mathrm{d}\mu(E)\frac{P_{n+1}(E)Q_{n}(-iz+E)-P_{n}(E)Q_{n+1}(-iz+E)}{h_{n}}~{}.$ (2.27)
@@
$z^{2}C(z)$ $=i\sum\limits_{n=0}^{\infty}\int\,\mathrm{d}\mu(E)\left\{\frac{2(\Delta_{n}-\Delta_{n+1})}{h_{n}}P_{n}(E)Q_{n}(-iz+E)\right.$ (2.28)
@@
If $\mathcal{K}$ is finite-dimensional, then the spectrum is discrete and finite, and $d_{\mathcal{K}}$ is an upper bound of the complexity $C(t)$. At late times, $C(t)$ may approach a constant value, $C(t)\to C_{\infty}<d_{\mathcal{K}}$, known as the plateau [21, 22, 39, 23, 32], or it may oscillate, even maximally between zero and $d_{\mathcal{K}}$. Therefore, the presence of a late-time plateau in physical systems, in contrast to unlimited growth, is a simple consequence of the fact that any physical system ultimately has a finite number of degrees of freedom. If $d_{\mathcal{K}}$ is huge, it will typically take an exponentially long time to reach the plateau. A late-time plateau with more or less pronounced oscillations also appears to occur in cases with unbounded discrete spectra [25].
Systems with a very large number of degrees of freedom can be effectively described in a large-$N$ (thermodynamic) limit for time scales that are short compared to the time needed to approach the plateau. The continuum approximation is a useful tool [31, 25, 32] to study such systems. In the large-$N$ limit, the spectrum becomes continuous and may be either bounded or unbounded. The following general statements can be made [37, 38]. The spectrum is bounded if and only if both sets of Lanczos coefficients, $\Delta_{n}$ and $a_{n}$, are bounded. If, for $n\to\infty$, these coefficients approach limits, $\Delta_{n}\to\bar{\Delta}$ and $a_{n}\to\bar{a}$, then the spectrum is bounded with at most countably many points outside the interval $[\bar{a}-2\sqrt{\bar{\Delta}},\bar{a}+2\sqrt{\bar{\Delta}}]$, and the limits of this interval are limit points of the spectrum. The MP distribution is an example of such a spectrum and will be discussed in section 3.
@@ 2.4 Linear transformation of energy
$C^{\prime}(t)=C(|\alpha|t)~{},\qquad C^{\prime}(z)=\frac{1}{|\alpha|}C\left(\frac{z}{|\alpha|}\right)~{}.$ (2.37)
@@ 3.1 MP distribution from Lanczos algorithm
Consider the resolvent $Q_{0}(z)$, which can be calculated with the help of (2.21). More precisely, because $a_{n}$ and $\Delta_{n}$ are constants for $n>0$, one has $R_{n}(z)=R(z)$, which satisfies
@@
The sign is determined by the condition $Q_{0}(z)\to 0$ for $|z|\to\infty$. The spectral density that follows from (2.19) simply is the MP distribution (1.4).
@@ 3.2 Late-time state complexity
In this subsection, we shall calculate the late-time linear growth coefficient of this state complexity. Recall that $\lambda$ formally parameterizes energy, but one should keep in mind the possibility of a linear transformation, as discussed in subsection 2.4. The late-time behaviour of $C(t)$ is encoded in the small-$z$ expansion of the Laplace transform $C(z)$. In particular, a leading order behaviour $C(z)\sim\frac{C_{0}}{z^{2}}$ implies the late-time complexity $C(t)\sim C_{0}t$. This is relevant in the present case.
@@
$z^{2}C(z)=i\int\,\mathrm{d}\mu(\lambda)\left\{-2cQ_{0}(-iz+\lambda)-c\left[Q_{1}(-iz+\lambda)+P_{1}(\lambda)Q_{0}(-iz+\lambda)\right]\right\}~{}.$ (3.7)
@@
$z^{2}C(z)=ic\left[1+\int\,\mathrm{d}\mu(\lambda)(iz-2\lambda)Q_{0}(-iz+\lambda)\right]~{}.$ (3.8)
@@
$z^{2}C(z)=-ic\int\,\mathrm{d}\mu(\lambda)\int\,\mathrm{d}\mu(\lambda^{\prime})\frac{\lambda+\lambda^{\prime}}{-iz+\lambda-\lambda^{\prime}}~{}.$ (3.9)
@@
$z^{2}C(z)=-ic\left[A(1-A)\int\,\mathrm{d}\tilde{\mu}(\lambda)\frac{2iz\lambda}{z^{2}+\lambda^{2}}+(1-A)^{2}\int\,\mathrm{d}\tilde{\mu}(\lambda)\int\,\mathrm{d}\tilde{\mu}(\lambda^{\prime})\frac{\lambda+\lambda^{\prime}}{-iz+\lambda-\lambda^{\prime}}\right]~{}.$ (3.11)
@@
$C_{0}=\lim\limits_{z\to 0}z^{2}C(z)$ $=2\pi c(1-A)^{2}\int\limits_{\lambda_{-}}^{\lambda_{+}}\,\mathrm{d}\lambda\left(\frac{\,\mathrm{d}\tilde{\mu}}{\,\mathrm{d}\lambda}\right)^{2}\lambda$

### v3

@@ 1 Introduction
with $\lambda_{\pm}=\left(1\pm\sqrt{c}\right)^{2}$. More precisely, $D(\lambda)=\Omega f(\lambda)$ with $c=\Omega\operatorname{e}^{-S}\in(0,\infty)$. The significance of the MP distribution stems from the MP theorem [5], which can be stated as follows. Let $X$ be an $M\times N$ matrix with independent and identically distributed real or complex entries. Then, in the large $N$ limit with $M/N\to c\in(0,\infty)$ fixed, the asymptotic eigenvalue distribution of the $M\times M$ matrix $A=\frac{1}{N}XX^{\dagger}$ (also called the Wishart matrix) is universal and given by the MP distribution. See [6] for an example in the context of random matrix theory. The MP distribution also appears under the name of free Poisson distribution in Free Probability Theory [7]. Since MP’s original proof from 1967, many different proofs have appeared in the mathematical literature (see [8, 9, 10] and references therein). For physicists, a proof using Feynman diagrams [11] seems especially interesting.
@@ 2.3 State complexity
$\tilde{C}(z)=\int\limits_{0}^{\infty}\,\mathrm{d}t\,C(t)\operatorname{e}^{-zt}\qquad(\Re z>0)~{}.$ (2.24)
@@
$\tilde{C}(z)=\sum\limits_{n=1}^{\infty}n\int\,\mathrm{d}\mu(E)\int\,\mathrm{d}\mu(E^{\prime})\frac{P_{n}(E)P_{n}(E^{\prime})}{h_{n}[z+i(E-E^{\prime})]}~{}.$ (2.25)
@@
$\tilde{C}(z)=-i\sum\limits_{n=1}^{\infty}n\int\,\mathrm{d}\mu(E)\frac{P_{n}(E)Q_{n}(-iz+E)}{h_{n}}~{}.$ (2.26)
@@
$z\tilde{C}(z)=\sum\limits_{n=0}^{\infty}\int\,\mathrm{d}\mu(E)\frac{P_{n+1}(E)Q_{n}(-iz+E)-P_{n}(E)Q_{n+1}(-iz+E)}{h_{n}}~{}.$ (2.27)
@@
$z^{2}\tilde{C}(z)$ $=i\sum\limits_{n=0}^{\infty}\int\,\mathrm{d}\mu(E)\left\{\frac{2(\Delta_{n}-\Delta_{n+1})}{h_{n}}P_{n}(E)Q_{n}(-iz+E)\right.$ (2.28)
@@
If $\mathcal{K}$ is finite-dimensional, then the spectrum is discrete and finite, and $d_{\mathcal{K}}$ is an upper bound of the complexity $C(t)$. At late times, $C(t)$ may approach a constant value, $C(t)\to C_{\infty}<d_{\mathcal{K}}$, known as the plateau [40, 41], or it may oscillate, even maximally between zero and $d_{\mathcal{K}}$. Therefore, the presence of a late-time plateau in physical systems, in contrast to unlimited growth, is a simple consequence of the fact that any physical system ultimately has a finite number of degrees of freedom. If $d_{\mathcal{K}}$ is huge, it will typically take an exponentially long time to reach the plateau. A late-time plateau with more or less pronounced oscillations also appears to occur in cases with unbounded discrete spectra [26].
Systems with a very large number of degrees of freedom can be effectively described in a large-$N$ (thermodynamic) limit for time scales that are short compared to the time needed to approach the plateau. The continuum approximation is a useful tool [40, 32, 26, 33] to study such systems. In the large-$N$ limit, the spectrum becomes continuous and may be either bounded or unbounded. The following general statements can be made [38, 39]. The spectrum is bounded if and only if both sets of Lanczos coefficients, $\Delta_{n}$ and $a_{n}$, are bounded. If, for $n\to\infty$, these coefficients approach limits, $\Delta_{n}\to\bar{\Delta}$ and $a_{n}\to\bar{a}$, then the spectrum is bounded with at most countably many points outside the interval $[\bar{a}-2\sqrt{\bar{\Delta}},\bar{a}+2\sqrt{\bar{\Delta}}]$, and the limits of this interval are limit points of the spectrum. The MP distribution is an example of such a spectrum and will be discussed in section 3.
@@ 2.4 Linear transformation of energy
$C^{\prime}(t)=C(|\alpha|t)~{},\qquad\tilde{C}^{\prime}(z)=\frac{1}{|\alpha|}\tilde{C}\left(\frac{z}{|\alpha|}\right)~{}.$ (2.37)
@@ 3.1 MP distribution from Lanczos algorithm
Consider the resolvent $Q_{0}(z)$, which can be calculated with the help of (2.21). More precisely, with (3.1) one has $R_{n}(z)=R(z)$ for $n>0$, which satisfies
@@
The sign is determined by the condition $Q_{0}(z)\to 0$ for $|z|\to\infty$. The spectral density that follows from (2.19) is the MP distribution (1.4).
@@ 3.2 Late-time state complexity
In this subsection, we shall calculate the late-time linear growth coefficient of this state complexity. Recall that $\lambda$ formally parameterizes energy, but one should keep in mind the possibility of a linear transformation, as discussed in subsection 2.4. The late-time behaviour of $C(t)$ is encoded in the small-$z$ expansion of the Laplace transform $\tilde{C}(z)$. In particular, a leading order behaviour $\tilde{C}(z)\sim\frac{C_{0}}{z^{2}}$ implies the late-time complexity $C(t)\sim C_{0}t$. This is relevant in the present case.
@@
$z^{2}\tilde{C}(z)=i\int\,\mathrm{d}\mu(\lambda)\left\{-2cQ_{0}(-iz+\lambda)-c\left[Q_{1}(-iz+\lambda)+P_{1}(\lambda)Q_{0}(-iz+\lambda)\right]\right\}~{}.$ (3.7)
@@
$z^{2}\tilde{C}(z)=ic\left[1+\int\,\mathrm{d}\mu(\lambda)(iz-2\lambda)Q_{0}(-iz+\lambda)\right]~{}.$ (3.8)
@@
$z^{2}\tilde{C}(z)=-ic\int\,\mathrm{d}\mu(\lambda)\int\,\mathrm{d}\mu(\lambda^{\prime})\frac{\lambda+\lambda^{\prime}}{-iz+\lambda-\lambda^{\prime}}~{}.$ (3.9)
@@
$z^{2}\tilde{C}(z)=-ic\left[A(1-A)\int\,\mathrm{d}\tilde{\mu}(\lambda)\frac{2iz\lambda}{z^{2}+\lambda^{2}}+(1-A)^{2}\int\,\mathrm{d}\tilde{\mu}(\lambda)\int\,\mathrm{d}\tilde{\mu}(\lambda^{\prime})\frac{\lambda+\lambda^{\prime}}{-iz+\lambda-\lambda^{\prime}}\right]~{}.$ (3.11)
@@
$C_{0}=\lim\limits_{z\to 0}z^{2}\tilde{C}(z)$ $=2\pi c(1-A)^{2}\int\limits_{\lambda_{-}}^{\lambda_{+}}\,\mathrm{d}\lambda\left(\frac{\,\mathrm{d}\tilde{\mu}}{\,\mathrm{d}\lambda}\right)^{2}\lambda$